OpenStudy (anonymous):

2x + y = 3 x - 2y = -1 If equation two is multiplied by -2 and then the equations are added, the result is _____.

OpenStudy (science0229):

If we multiply equation 2 by -2, it would be -2x+4y=2.

OpenStudy (anonymous):

would it be like 3y=5?

OpenStudy (science0229):

2x+y=3 -2x+4y=2 5y=5 So y=1

OpenStudy (science0229):

Plug that in for either one of the equation and find x

OpenStudy (anonymous):

all i needed to know was that 5y=5